OpenStudy (anonymous):

Simplify: x^2-16/x^2-8x+16

OpenStudy (anonymous):

:/ Maybe this can help? http://www.wolframalpha.com/input/?i=%28x%5E2-16%29%2F%28x%5E2-8x%2B16%29

OpenStudy (anonymous):

alright, i checked my answer with it & i got it right

OpenStudy (anonymous):

ok let me do this since \[\Large (a^2-b^2)=(a+b)(a-b)\] then \[\Large (x^2-16)=(x^2-(4)^2)=(x+4)(x-4)\] now factoring the denominator

OpenStudy (lgbasallote):

\[\large \frac{x^2 - 16}{x^2 -8x + 16} \implies \frac{(x+4)(x-4)}{(x-4)(x-4)} \implies \frac{x+4}{x-4}\]

OpenStudy (anonymous):

yes, thats what i got

OpenStudy (anonymous):

\[\Large (x^2-8x-16)=(x^2-4x-4x+16)\] \[\Large (x^2-4x-4x+16)=(x(x-4)-4(x-4))=(x-4)(x-4)\]